| Name | Email | Date | Comment
|
| ANS
| anolla@naic.edu
|
10/7/99
| A line filter was added to the supply cryo line.
|
| ANS
| anolla@naic.edu
|
9/28/2000
| Receiver was not working , due that the AC power for the Lin/Pol box was disconected.
|
| ANS
| anolla@naic.edu
|
10/4/2000
| Again the Lin/Pol box was not working, this time was found that the control cable for the
box was broken at the rfi filters on the cabinet.
|
| ANS
| anolla@naic.edu
|
10/25/2000
| System shut off for maint.Crosshead higly contaminated ,the crosshead ED 027
was replaced with ED 021 , but after installing these replacement crosshead it got bad for
some reason ,so cleaning up ED 027 to installed it back.
|
| ANS
| anolla@naic.edu
|
10/27/2000
| Problems with crosshead ED027 , it was leaking thru the power conn. It was
replaced with ED 021.
|
| ANS
| anolla@naic.edu
|
11/25/2000
| Cryo comp 011 blew up, so cryo system hooked to comp 007.
|
| ANS
| anolla@naic.edu
|
12/27/2000
| Cryo system hooked up to comp 012 , was using comp 007.
|
| ANS
| anolla@naic.edu
|
12/29/2000
| Cryo system hooked back to comp 007 , but using cryo lines 3 & 4. We suspect of
leaky lines on 12 & 13.
|
| ANS
| anolla@naic.edu
|
1/19/2001
| The Vacion pump power supply turned off , it was causing RFI.It happened with
both power supplies at the Dome.
|
| ANS
| anolla@naic.edu
|
` 3/12/2001
| The cryo comp failed on 3/10/2001 due to a power glitch.Jose Rosa & Bill Genter
came to cooled it down on 3/11/2001.
|
| ANS
| anolla@naic.edu
|
7/5/2001
| The polarization sw was not working due to a bad 5 vdc supply (shorted out
capacitor on the regulator)
|
| ANS
| anolla@naic.edu
|
7/16/2001
| The polarization sw box was taken down on 7/13/2001 in order to check its phase
shift across it.
|
| ANS
| anolla@naic.edu
|
7/18/2001
| Polarization box installed back into system by EC/EG . It was taken down to
measure its phase shift because it was not in schedule its use.
|
| ANS
| anolla@naic.edu
|
8/13/2001
| The cryo comp and cryo lines were hooked to their proper ones. Now comp ED
008 and lines 8/9 are the ones hooked to the system as they were.
|
| ANS
| anolla@naic.edu
|
8/21/2001
| Due to a power glitch the cryo comp tripped off. It was recovered after 15
minutes.
|
| ANS
| anolla@naic.edu
|
10/2/2001
| Cryo compressor failed ( ED 008) , sound like broken flappers inside of it . It was
replaced with comp ED 011.
|
| ANS
| anolla@naic.edu
|
10/9/2001
| System off in order to replace crosshead, is overdue.
|
| ANS
| anolla@naic.edu
|
10/10/2001
| System was cooled down without replacing the crosshead due to running short of
time, the 430 mhz dual-beam experiment was running and we were no able to stay
upstairs. Have to be reschedule again, crosshead replacement soon overdue.
|
| ANS
| anolla@naic.edu
|
10/17/2001
| The system was shutted off on 10/15/2001 in order to replace the crosshead. The
crosshead removed was ED 021 and the one installed ED 024.
|
| ANS
| anolla@naic.edu
|
10/18/2001
| The amps were turned on today . Cryo comp was 50/245 , charged it to 65/260.
|
| ANS
| anolla@naic.edu
|
2/8/2002
| Due that cryo temp was increasing Coco went upstairs and purged the system; it is
on schedule to replace the cryopump next week.
|
| ANS
| anolla@naic.edu
|
2/11/2002
| Replacement RF filters for the filter bank pos # 6 arrived, the ones that were there
had a high ripple on them. These was done on 11/15/2001 , there is no entry of such
problem on this logbook. Fo is 1415 mhz B.W. 110 mhz. (1360/1470 mhz).
|
| ANS
| anolla@naic.edu
|
2/14/2002
| System shut off in order to replace the cryopump , it is contaminated . Also the
polarization box and filter bank was taken down to check them.
|
| ANS
| anolla@naic.edu
|
2/15/2002
| Cryopump replaced was ED 024 , now is ED 027 (1020's). System been cooled
down. The polarization box and filter bank were removed in order to check them down at
the lab.
|
| ANS
| anolla@naic.edu
|
2/18/2002
| System back to normal , the pol box and filter bank were installed back.
|
| ANS
| anolla@naic.edu
|
2/20/2002
| JR found the cryopump loose (system shut off), it had the back holding screws
broken . Were removed using an screw extractor , also treads of screws were cleaned.By
mistake we left Dewar pumping and after 24 hours the Dewar was still cold , so when we
took out the cryopump it's displacers got frozen ,beware next time,don't leave Dewar
pumping.
|
| ANS
| anolla@naic.edu
|
2/21/2002
| System been cooled ,using new tunable cryopump driver. Still the squique noise
can be heared on the cryopump.
|
| ANS
| anolla@naic.edu
|
2/22/2002
| The Dewar vacuum was found reading to high this morning ( 6 un's) , also VP
was high ( 50 ) and its cryo temp were 26.7 ºk at 16 º k stage and 165.7 ºk at 70 ºk stage
, then we started pumping the Dewar and its cryo parameters went down ( vac 1.2 un's ,
VP 9 , 16 º stage 18.9 ºk , 70 ºk stage 158.7 ºk ) . There is also a problem with the
downstairs cryo reading , it is not accurate have to find out what's wrong with it.
Went up again in the afternoon to double check the Dewar status . Its cryo temps
were 18.1 at 16 º stage and 140.1 at the 70 º stage , VP was 7 and vacuum 1.8 un's .
|
| ANS
| anolla@naic.edu
|
2/25/2002
| Found this morning that cryo temp , vp & vac were high , so we start pumping it .
The vac valve at the Dewar left open with the pump running.
|
| ANS
| anolla@naic.edu
|
2/26/2002
| The Dewar vac was 1 un and its VP 2, also its cryo temps were 15.7 & 104.9 after
been pumping it for 24 hours. So we are monitoring it for a while. Valve was closed.
|
| ANS
| anolla@naic.edu
|
2/28/2002
| JR brought down the rem / local cryo reading chassis to check it due that the
remote reading was giving wrong readings. IC found a loose connection on the relay
board and replace the relay. Now the readings are ok.
|
| ANS
| anolla@naic.edu
|
3/7/2002
| JR found Dewar was losing vacuum , was 2.5 un's.VP was 9 . Cryo temps were
18.9 ºk and 114.7 ºk . So he started pumping the Dewar we suspect of a small leak or He
contamination .
|
| ANS
| anolla@naic.edu
|
3/8/2002
| Valve closed , VP 1 , vac 1.4 un's . Cryo temps are 14.7 ºk and 89.9 ºk.
|
| ANS
| anolla@naic.edu
|
4/2/2002
| System turned off in order to replace cryopump, cryo temp has been cycling since
last cool down.
|
| ANS
| anolla@naic.edu
|
4/3/2002
| Replace the cryopump , it was 027 , now is 024.
|
| ANS
| anolla@naic.edu
|
4/15/2002
| The recvr Dewar has a leak since ??????? , so the vacpump was left pumping the
Dewar. It has a better cryo temps .WG & JR were triying to locate it with the Varian leak
detector.
|
| ANS
| anolla@naic.edu
|
4/17/2002
| System still with a leak. Vacpump attached to Dewar again.
|
| ANS
| anolla@naic.edu
|
6/5/2002
| Vacpump attached to Dewar , vpg & cryotemps up again.
|
| ANS
| anolla@naic.edu
|
6/7/2002
| Dewar valve closed.Vac pump off.
|
| ANS
| anolla@naic.edu
|
6/13/2002
| Vacpump attached to Dewar, VPG was 8.Its vacion current is typically 80 ua.
|
| ANS
| anolla@naic.edu
|
8/1/2002
| WG swap the bias card in order to pinpoint an oscillation in channel
A.Biases were adjusted, Tsys normal in both channels after the swap.
Biases settings were:Stage 1: 1.538V/7.52 mA Stage 2: 2.518V/7.04 mA Stage 3: 3.903/8.04 mA Chann B
|
| ANS | anolla@naic.edu | Wednesday, September 25, 2002 11:51:28 | Dewar warmed up due to power dip , been cooled.
|
| ans | anolla@naic.edu | Thursday, September 26, 2002 14:13:10 | Dewar cooled. Dewar valve closed . Vacion on ( 70 ua )
|
| ans | anolla@naic.edu | Wednesday, October 9, 2002 12:51:27 | The cryo comp was relpaced with one with remote sw (ED 010)it was cryocomp ED 013.
|
| ans | anolla@naic.edu | Monday, October 28, 2002 09:51:52 | Cryocomp replaced due to fan failure.
|
| ans | anolla@naic.edu | Thursday, November 21, 2002 11:23:07 | The receiver focal point was readjusted . It was lowered by 4".Have to observe with it to know if there is any improment.
|
| LSW | lwray@naic.edu | Friday, November 22, 2002 17:41:23 | improvement due to receiver re-focussing: 10% increase in K/Jy of the telescope.
|
| ans | anolla@naic.edu | Wednesday, November 13, 2002 13:57:42 | Amp sn L126 is on channel A , sn L128 is on B. Also the power barrier strips were isolated with insulators.
|
| ans | anolla@naic edu. | Tuesday, January 28, 2003 11:28:09 | Control boxes were removed from system , going to be tested with the new NRAO recvr.
|
| ans | anolla@naic.edu | Monday, February 3, 2003 13:30:40 | System was shut off in order to remove the Dewar to replce it with the NRAO one.
|
| ans | anolla@naic edu. | Tuesday, February 18, 2003 14:32:01 | Dewar removed & taken down.
|